# Paper/main_online_appendix

### v3

@@ B.6 Intertemporal Preference Reversals: Details
$TE(c,Z)$ =$=\begin{cases}1/2[t_{R(c,Z)-1}+t_{R(c,Z)}]&R(c,Z)<n+1\\ t_{n}+1/2(t_{n}-t_{n-1})&R(x,Z)=n+1\end{cases}$
@@ B.8 Relationship to Random Utility
Suppose $\rho$ has an $L_{1}$-complexity representation $(\beta,G)$, where $G(1)=G(-1)=1$. Then, there exists an $n$-dimensional random vector $\tilde{\beta}$ such that for all $(x,y)\in\mathcal{D}$, $\rho(x,y)=\mathbb{P}\left\{\sum_{k}\tilde{\beta}_{k}x_{k}\geq\sum_{k}\tilde{\beta}_{k}y_{k}\right\}$, where for all $k$, $\mathbb{P}\left\{\text{sgn}(\tilde{\beta}_{k})=\text{sgn}(\beta_{k})\right\}=1$.
@@ B.9 Relationship to Linear Differentiation Model
In this section, we contrast the $L_{1}$-Complexity model (Definition 2) against the Linear Differentiation Model (LDM) proposed in He and Natenzon, 2023b .
@@
(He and Natenzon, 2023b, ). A binary choice rule $\rho$ has a linear differentiation representation if there exists $\beta\in\mathbb{R}^{n}$, an $n\times n$ symmetric positive-definite matrix $\Sigma$, and a continuous, strictly increasing function $G$, such that
@@
The intuition for the monotonicity violations produced by the LDM stem stems from a property of the model formalized in Proposition 1 from He and Natenzon, 2023b , which states that any $x$ that solves $\max_{x}\rho(x,y)$ lies along along the same one-dimensional subset of $\mathbb{R}^{n}$ containing $y$. An immediate implication of this property is that for any $x^{\prime}$ that dominates $x$, we will have $\rho(x^{\prime},y)<\rho(x,y)$ so long as $x^{\prime}$ does not also lie in that subset.
@@ B.10 Relationship to Bayesian Probit
00 0 $\leq 1+2\sigma_{xy}\sigma_{xz}\sigma_{yz}-\sigma_{xy}^{2}\sigma_{yz}^{2}\sigma_{xz}^{2}.$
@@ C.1 Characterization of $L_{1}$-Complexity
Call a (complete) binary relation $\succeq$ on $X$ the stochastic order induced by a binary choice rule $\rho$ if for all $x\neq y$, $x\succeq y$ if $\rho(x,y)\geq 1/2$, and for all $x\in X$, $x\succeq x$. Say that a binary choice rule $\rho$ satisfies moderate transitivity if for $\rho(x,y),\rho(y,z)\geq 1/2$, then $\rho(x,z)>\min\{\rho(x,y),\rho(y,z)\}$ or $\rho(x,z)=\rho(x,y)=\rho(y,z)$. Say that a binary choice rule $\rho$ satisfies weak transitivity if for $\rho(x,y),\rho(y,z)\geq 1/2$, $\rho(x,y)\geq 1/2$. Consider a partial order $\geq_{X}$ on $X$. Say that a binary choice rule $\rho$ satisfies monotonicity with respect to $\geq_{X}$ if $x^{\prime}\geq_{X}x$ implies $\rho(x^{\prime},y)\geq\rho(x,y)$, where the inequality is strict whenever $x\not\geq_{X}x^{\prime}$, $x\not\geq_{X}y$ and $y\not\geq_{X}x$. Say that $\rho$ satisfies dominance with respect $\geq_{X}$ if whenever $x\geq_{X}y$, we have $\rho(x,y)\geq\rho(w,z)$ for all $w,z\in X$, where the inequality is strict if $w\not\geq_{X}z$.
@@
For $z\in\mathbb{R}^{n}$, Let $d^{+}(z)=\sum_{k:z_{k}\geq 0}z_{k}$ and $d^{-}(x)=\sum_{k:z_{k}<0}|z_{k}|$ denote the summed advantages and disadvantages in the comparison between $z$ and 00. Say that $z$ has no dominance relationship if $d^{+}(z),d^{-}(z)>0$.
@@
For $i,j\in\{1,...,n\}$, $i\neq j$, define $z^{ij}\in\mathbb{R}^{n}$ satisfying
@@
Take any $x,y$, and let $z=x-y$. Due to linearity, we have
@@
$=G\left(\frac{U(x)-U(y)}{d_{L1}(x,y)}\right)$
as desired.
@@ C.2 Multinomial Choice Results
Consider a continuous distribution with density $g$ that is symmetric around 00 and satisfies the monotone likelihood ratio property. The function $g$ then has the following properties:
@@
So for any $y>x$, $g^{\prime}(y)>0$. Symmetry implies that for any $y<-x$, $g^{\prime}(y)<0$, and so $g$ is not integrable, a contradiction. To see 3), note that by symmetry, $\frac{g(t-x)}{g(-t-x)}=1$ for $x=0$. MLRP of $g$ implies that $\frac{g(t-x)}{g(-t-x)}>1$ for all $x>0$ as desired. ∎
The following observations pertain to finite set of options $A$. Enumerate $A$ by $1,2,...,N$ and let $s=(s_{ij})_{i<j}$ collect all pairwise signals in $A$. Let $X_{(k)}^{N}$ denote the $k$th order statistic among $N$ draws from the prior distribution $q$. Let $V_{(k)}^{N}=\mathbb{E}[X_{(N+1-k)}]$, that is, $V_{(k)}^{N}$ gives the expected value of an option if it is ranked $k$th. Let $\pi:A\to A$ denote a permutation function; let $\Pi$ denote the set of permuation functions on $A$. With some abuse of notation, associate each $\pi$ with the event that the $v_{i}$’s are ordered according to $\pi$: that is $\pi(i)=n$ means that option $i$ is ranked $n$th in the ordering. The posterior expected value of an option $i$ given signal $s$ is then given by
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
To show (i), consider the case where $\tau=0$. We have that with probability 1, $p_{k}(s)=1/{n+1}$ for all $k\in\{1,...,n+1\}$. Let $\mu$ denote the expectation of $Q$. By symmetry of $Q$, we have $V^{N}_{(k)}=2\mu-V^{N}_{(N+1-k)}$ for all $k=1,...,N$, and so $\mathbb{E}[v_{x}|s]=\mu$ with probability 1.

### v4

@@ B.6 Intertemporal Preference Reversals: Details
$TE(c,Z)$ $=\begin{cases}1/2[t_{R(c,Z)-1}+t_{R(c,Z)}]&R(c,Z)<n+1\\ t_{n}+1/2(t_{n}-t_{n-1})&R(x,Z)=n+1\end{cases}$
@@ B.8 Relationship to Random Utility
Suppose $\rho$ has an $L_{1}$-complexity representation $(\beta,G)$, where G(1)=G(1)=1G(1)=G(-1)=1. Then, there exists an $n$-dimensional random vector $\tilde{\beta}$ such that for all $(x,y)\in\mathcal{D}$, $\rho(x,y)=\mathbb{P}\left\{\sum_{k}\tilde{\beta}_{k}x_{k}\geq\sum_{k}\tilde{\beta}_{k}y_{k}\right\}$, where for all $k$, $\mathbb{P}\left\{\text{sgn}(\tilde{\beta}_{k})=\text{sgn}(\beta_{k})\right\}=1$.
@@ B.9 Relationship to Linear Differentiation Model
In this section, we contrast the $L_{1}$-Complexity model (Definition 2) against the Linear Differentiation Model (LDM) proposed in He and Natenzon, (2023).
@@
(He and Natenzon,, 2023). A binary choice rule $\rho$ has a linear differentiation representation if there exists $\beta\in\mathbb{R}^{n}$, an $n\times n$ symmetric positive-definite matrix $\Sigma$, and a continuous, strictly increasing function $G$, such that
@@
The intuition for the monotonicity violations produced by the LDM stem stems from a property of the model formalized in Proposition 1 from He and Natenzon, (2023), which states that any $x$ that solves $\max_{x}\rho(x,y)$ lies along along the same one-dimensional subset of $\mathbb{R}^{n}$ containing $y$. An immediate implication of this property is that for any $x^{\prime}$ that dominates $x$, we will have $\rho(x^{\prime},y)<\rho(x,y)$ so long as $x^{\prime}$ does not also lie in that subset.
@@ B.10 Relationship to Bayesian Probit
$0$ $\leq 1+2\sigma_{xy}\sigma_{xz}\sigma_{yz}-\sigma_{xy}^{2}\sigma_{yz}^{2}\sigma_{xz}^{2}.$
@@ C.1 Characterization of $L_{1}$-Complexity
Call a (complete) binary relation $\succeq$ on $X$ the stochastic order induced by a binary choice rule $\rho$ if for all $x\neq y$, $x\succeq y$ if $\rho(x,y)\geq 1/2$, and for all $\in X, xx$$\succeq x$. Say that a binary choice rule $\rho$ satisfies moderate transitivity if for $\rho(x,y),\rho(y,z)\geq 1/2$, then $\rho(x,z)>\min\{\rho(x,y),\rho(y,z)\}$ or $\rho(x,z)=\rho(x,y)=\rho(y,z)$. Say that a binary choice rule $\rho$ satisfies weak transitivity if for $\rho(x,y),\rho(y,z)\geq 1/2$, $\rho(x,y)\geq 1/2$. Consider a partial order $\geq_{X}$ on $X$. Say that a binary choice rule $\rho$ satisfies monotonicity with respect to $\geq_{X}$ if $x^{\prime}\geq_{X}x$ implies $\rho(x^{\prime},y)\geq\rho(x,y)$, where the inequality is strict whenever $x\not\geq_{X}x^{\prime}$, $x\not\geq_{X}y$ and $y\not\geq_{X}x$. Say that $\rho$ satisfies dominance with respect $\geq_{X}$ if whenever $x\geq_{X}y$, we have $\rho(x,y)\geq\rho(w,z)$ for all $w,z\in X$, where the inequality is strict if $w\not\geq_{X}z$.
@@
For $z\in\mathbb{R}^{n}$, Let $d^{+}(z)=\sum_{k:z_{k}\geq 0}z_{k}$ and $d^{-}(x)=\sum_{k:z_{k}<0}|z_{k}|$ denote the summed advantages and disadvantages in the comparison between $z$ and 0. Say that $z$ has no dominance relationship if $d^{+}(z),d^{-}(z)>0$.
@@
For $\in\{1,...,n\}, iji\neq j$, define $z^{ij}\in\mathbb{R}^{n}$ satisfying
@@
Take any $x,y$, and let z=xyz=x-y. Due to linearity, we have, as desired
@@
$=G\left(\frac{U(x)-U(y)}{d_{L1}(x,y)}\right).$
@@ C.2 Multinomial Choice Results
Consider a continuous distribution with density $g$ that is symmetric around 0 and satisfies the monotone likelihood ratio property. The function $g$ then has the following properties:
@@
So for any $y>x$, $g^{\prime}(y)>0$. Symmetry implies that for any y<xy<-x, $g^{\prime}(y)<0$, and so $g$ is not integrable, a contradiction. To see 3), note that by symmetry, $\frac{g(t-x)}{g(-t-x)}=1$ for $x=0$. MLRP of $g$ implies that $\frac{g(t-x)}{g(-t-x)}>1$ for all $x>0$ as desired. ∎
The following observations pertain to finite set of options $A$. Enumerate $A$ by 1,2,,N1,2,...,N and let $s=(s_{ij})_{i<j}$ collect all pairwise signals in $A$. Let $X_{(k)}^{N}$ denote the $k$th order statistic among $N$ draws from the prior distribution $q$. Let $V_{(k)}^{N}=\mathbb{E}[X_{(N+1-k)}]$, that is, $V_{(k)}^{N}$ gives the expected value of an option if it is ranked $k$th. Let $\pi:A\to A$ denote a permutation function; let $\Pi$ denote the set of permuation functions on $A$. With some abuse of notation, associate each $\pi$ with the event that the $v_{i}$’s are ordered according to $\pi$: that is $\pi(i)=n$ means that option $i$ is ranked $n$th in the ordering. The posterior expected value of an option $i$ given signal $s$ is then given by
@@
$\rho(y;x|z)=$ $\int_{e_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e_{yz}=-\infty}^{\infty}1\left\{\mathbb{E}[v_{y}|s]-\mathbb{E}[v_{x}|s]\geq 0\right\}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
$+\int_{e_{xz}=\sqrt{\tau_{xz}}}^{\infty}\int_{e_{yz}=-\infty}^{\infty}1\left\{\mathbb{E}[v_{y}|s]-\mathbb{E}[v_{x}|s]\geq 0\right\}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
@@
$\int_{e_{xz}=\sqrt{\tau_{xz}}}^{\infty}\int_{e_{yz}=-\infty}^{\infty}1\left\{\mathbb{E}[v_{y}|s]-\mathbb{E}[v_{x}|s]\geq 0\right\}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
$=\int_{e_{xz}=\sqrt{\tau_{xz}}}^{\infty}\int_{e_{yz}=-\infty}^{\infty}1\left\{g(e_{xz}-2\sqrt{\tau_{xz}})g(e_{yz})-g(e_{xz})g(e_{yz}-2\sqrt{\tau_{yz}})\geq 0\right\}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
$=\int_{e^{\prime}_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e^{\prime}_{yz}=-\infty}^{\infty}1\left\{g(e^{\prime}_{xz})g(e^{\prime}_{yz}-2\sqrt{\tau_{yz}})-g(e^{\prime}_{xz}-2\sqrt{\tau_{xz}})g(e^{\prime}_{yz})\geq 0\right\}g(e^{\prime}_{xz}-2\sqrt{\tau_{xz}})g(e^{\prime}_{yz}-2\sqrt{\tau_{yz}})de^{\prime}_{yz}de^{\prime}_{xz}$
$=\int_{e^{\prime}_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e^{\prime}_{yz}=-\infty}^{\infty}1\left\{\mathbb{E}[v_{y}|s]-\mathbb{E}[v_{x}|s]\leq 0\right\}g(e^{\prime}_{xz}-2\sqrt{\tau_{xz}})g(e^{\prime}_{yz}-2\sqrt{\tau_{yz}})de^{\prime}_{yz}de^{\prime}_{xz}$
@@
$\rho(y;x|z)=$ $\int_{e_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e_{yz}=e^{*}_{yz}(\tau_{xz},e_{xz})}^{\infty}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
$+\int_{e^{\prime}_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e_{yz}=-\infty}^{e^{*}_{yz}(\tau_{xz},e_{xz})}g(e_{xz}-2\sqrt{\tau_{xz}})g(e_{yz}-2\sqrt{\tau_{yz}})\,de_{yz}de_{xz}$
$=$ $\int_{e_{xz}=-\infty}^{\sqrt{\tau_{xz}}}\int_{e_{yz}=e^{*}_{yz}(\tau_{xz},e_{xz})}^{\infty}g(e_{xz})g(e_{yz})\,de_{yz}de_{xz}$
$+\int_{e_{xz}=-\infty}^{-\sqrt{\tau_{xz}}}\int_{e_{yz}=-\infty}^{e^{*}_{yz}(\tau_{xz},e_{xz}+2\sqrt{\tau_{xz}})}g(e_{xz})g(e_{yz}-2\sqrt{\tau_{yz}})\,de_{yz}de_{xz}$
@@
$=\frac{1}{2\sqrt{\tau_{xz}}}\left[\int_{e_{yz}=e^{*}_{yz}\left(\tau_{xz},\sqrt{\tau_{xz}}\right)}^{\infty}g\left(\sqrt{\tau_{xz}}\right)g(e_{yz})\,de_{yz}-\int_{e_{yz}=-\infty}^{e^{*}_{yz}\left(\tau_{xz},\sqrt{\tau_{xz}}\right)}g\left(-\sqrt{\tau_{xz}}\right)g(e_{yz}-2\sqrt{\tau_{yz}})\,de_{yz}\right]$
@@
$\quad+\int_{e_{xz}=-\infty}^{-\sqrt{\tau_{xz}}}\left[\frac{\partial}{\partial\tau_{xz}}e^{*}_{yz}(\tau_{xz},e_{xz}+2\sqrt{\tau_{xz}})+\frac{1}{\sqrt{\tau_{xz}}}\frac{\partial}{\partial e_{xz}}e^{*}_{yz}(\tau_{xz},e_{xz}+2\sqrt{\tau_{xz}})\right]g(e_{xz})g(e^{*}_{yz}(\tau_{xz},e_{xz}+2\sqrt{\tau_{xz}})-2\sqrt{\tau_{yz}})\,de_{xz}$
@@
$\quad+\int_{e_{xz}=-\infty}^{-\sqrt{\tau_{xz}}}\frac{1}{\sqrt{\tau_{xz}}}\frac{\partial}{\partial e_{xz}}e^{*}_{yz}(\tau_{xz},e_{xz})g(e_{xz})g(e^{*}_{yz}(\tau_{xz},e_{xz}+2\sqrt{\tau_{xz}})-2\sqrt{\tau_{yz}})\,de_{xz}$
@@
To show (i), consider the case where $\tau=0$. We have that with probability 1, $p_{k}(s)=1/{n+1}$ for all $k\in\{1,...,n+1\}$. Let $\mu$ denote the expectation of $Q$. By symmetry of $Q$, we have $V^{N}_{(k)}=2\mu-V^{N}_{(N+1-k)}$ for all k=1,,Nk=1,...,N, and so $\mathbb{E}[v_{x}|s]=\mu$ with probability 1.